‘Darshan’ is an Internet-based travel agency wherein customers can see videos of the cities they plan to visit. The number of hits daily is a normally distributed random variable with a mean of 10,000 and a standard deviation of 2,400.

A. What is the probability of getting fewer than 9,000 hits? 


1
Expert's answer
2022-03-23T16:55:45-0400

"P(X<9000)=P(Z<\frac{9000-10000}{2400})=P(Z<-0.42)=0.3372."